### v2

# LightRetriever: A LLM-based Hybrid Retrieval Architecture with 1000× Faster Query Inference

Abstract
Large Language Models (LLMs)-based hybrid retrieval uses LLMs to encode queries and documents into low-dimensional dense or high-dimensional sparse vectors. It retrieves documents relevant to search queries based on vector similarities. Documents are pre-encoded offline, while queries arrive in real-time, necessitating an efficient online query encoder. Although LLMs significantly enhance retrieval capabilities, serving deeply parameterized LLMs slows down query inference throughput and increases demands for online deployment resources. In this paper, we propose LightRetriever, a novel LLM-based hybrid retriever with extremely lightweight query encoders. Our method retains a full-sized LLM for document encoding, but reduces the workload of query encoding to no more than an embedding lookup. Compared to serving a full-sized LLM on an H800 GPU, our approach achieves over a 1000× speedup for query inference with GPU acceleration, and even a 20× speedup without GPU. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse retrieval tasks, retaining an average of 95% full-sized performance.
1 Introduction
Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities [62]. The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval. Dense retrieval [29] compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity. Meanwhile, sparse retrieval [60] characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) [3, 20] and BM25 [60]. Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores [72] or ranks [15]. Text retrieval is broadly applied in various scenarios, such as web search [43], semantic textual similarity [8], fact checking [66], and retrieval augmented generation [39].
Figure 1: Architecture comparison of LLM-based hybrid retrieval with 1) Full-sized Query Inference (Previous works); and 2) Extreme Lightweight Query Inference (LightRetriever).
Deploying retrieval systems requires different processing workflows for queries and documents. As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases [27, 60]. Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere [79]. On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable. The query encoder $Enc_{q}$ needs to be served online. Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM. Most existing works of text retrieval [29, 11, 71] utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides. Recent state-of-the-art (SOTA) retrievers [71, 5] employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities on various languages and tasks. However, serving giant LLMs slows down query inference throughput and depletes online deployment resources [37]. Nowadays, LLM-based text retrieval incurs non-negligible online workloads, necessitating special optimizations for query inference efficiency.
Modern Transformers-based LMs handle text modeling with the attention mechanism [68], including the contextual traits and deep interactions of co-occurrences on discrete tokens within certain texts. As is the case for existing LLM-based text retrievers [71, 5], queries and documents are modeled equally with the same LLM, enduring the same computational burdens. In contrast, traditional statistical retrieval algorithms, such as the popular TF-IDF [63] and BM25 [60], only characterize limited token co-occurrence by analyzing lexical overlaps with customized retrieval metrics. They do not handle the contextual traits and deep co-occurrences, and their encoders do not need to be deeply parametrized during inference. Experiments from [65] show that BM25 still performs well in many retrieval tasks, even beating some previous BERT-based retrievers trained on MSMARCO [52].
Built on the above observations, to minimize query service cost, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture targeted at extreme lightweight query inference abilities. As is shown in Figure 1, we retain a full-sized LLM on the document side ($Enc_{d}$) for strong retrieval capacities, but completely remove the deep modeling of queries. 1) For dense retrieval, our method end-to-end trains cacheable token vectors from the LLM-based query encoder. The dense query representations are simply aggregated by averaging corresponding token vectors shown in the texts. All query token vectors are pre-computed and cached into a simple Embedding layer. Query inference costs are drastically reduced to only one Embedding lookup. 2) For sparse retrieval, similar to statistical retrieval algorithms, the mapping of a token and its corresponding frequency [^1] Here we directly use the number of a certain token that appeared in the tokenized query as its frequency. is directly used as the sparse query vector. Thus, no more encoding stage for sparse retrieval is needed. The dense similarity scores are computed with cosine similarity, while the sparse scores directly use inner product similarity. These scores are linearly interpolated [72] for final hybrid retrieval.
LightRetriever is super fast and lightweight with respect to query inference. Compared to serving a full-sized LLM (e.g., Mistral-7B [26]) on an H800 GPU, our method achieves over a 1000× speedup for query inference with GPU acceleration, and even a 20× speedup without GPU. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Mistral-7B, Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR [65] and 8 Chinese tasks from CMTEB Retrieval [76]), retaining an average of 95% full-sized performance.
2 Algorithm
Figure 2: Full-sized symmetric dense retrieval for baseline training.
2.1 LM-based text retrieval
Definition
Given a search query $q$ and candidate documents $d=\{d^{+};d^{-}\}\in\mathcal{N}(q)$ (including irrelevant documents $d^{-}$ and relevant documents $d^{+}$) retrieved from indexes $\mathcal{N}$, LM-based text retrieval aims to learn optimized encoders $Enc_{q}$ and $Enc_{d}$ with parameters $\theta$, where the similarity of the query vector $v_{q}=Enc_{q}(q)$ and irrelevant document vectors $v_{d^{-}}=Enc_{d}(d^{-})$ are encouraged to be smaller than those of relevant documents $v_{d^{+}}=Enc_{d}(d^{+})$. Assuming the dot product is used as the similarity metric, the above optimization goals can be described as the following objective,
$\min_{\theta}\max_{d^{-}\in\mathcal{N}(q)}\mathcal{E}(q,d^{+},d^{-};\theta)\text{, where }\mathcal{E}(q,d^{+},d^{-};\theta)=v_{q}\cdot v_{d^{-}}-v_{q}\cdot v_{d^{+}}$ (1)
which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ [12]. The above objective can be effectively optimized by the pairwise/listwise contrastive loss [10], margin loss [61], etc. Following [29], listwise contrastive loss ($\ell^{CL}$) is used in our work,
$\ell^{CL}=-\log\frac{e^{v_{q}\cdot v_{d^{+}}/\tau}}{e^{v_{q}\cdot v_{d^{+}}/\tau}+\sum_{d^{-}\in\mathcal{N}(q)}e^{v_{q}\cdot v_{d^{-}}/\tau}}.$ (2)
, where $\tau$ is the temperature. The retrieval capacity is strongly correlated with parametrized encoders. As discussed before, serving deep parametrized LMs, especially giant LLMs, consumes large amounts of deployment resources. Thus, our work aims to address this issue by designing a novel asymmetric retrieval architecture with extreme imbalance.
Instructed retrieval
Similar to instructed fine-tuning (SFT) [55], modern LM-based retrievers [64, 76, 70], especially LLM retrievers [71, 5, 38], customize task specific instructions for better domain adaption abilities. Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval. Although [64, 76] use instructions on both query and passage sides, most LLM-based retrievers [71, 5, 38] do not use passage instructions to reuse the pre-built indexes across different tasks. In our work, dense retrieval follows the query instructions of Mistral-E5 [71] and does not involve passage instructions. The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders. Detailed instructions are shown in the Appendix A.4.
2.2 Dense retrieval of LightRetriever
For LLM-based retrieval, LightRetriever pushes the query inference efficiency to a new extreme by completely removing the deep modeling on the query side. For dense retrieval, our work end-to-end trains cacheable token vectors for online serving.
Figure 3: Dense retrieval of LightRetriever.
Training
Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction. Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^2] Specifically, EOS pooling is used in our work by following Mistral-E5 [71]. with the query encoder [^3] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.5.,
$v_{t_{i}}^{\text{den}}=Enc_{q}(Inst;t_{i})$ (3)
Then, the dense query vector $v_{q}^{\text{den}}$ of the whole query text is aggregated by averaging token vectors,
$v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{\text{den}}$ (4)
Contrastive loss is the main objective for dense training, as defined in Equation 2.
Following [9, 32], an auxiliary KL loss is also used to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
$\ell^{Align}=\texttt{KL-Div}(S^{student},S^{teacher})$ (5)
Full-sized symmetric baselines could be trained with asymmetric encoders together. Thus, the computation of auxiliary loss is economical. Experiments show that this helps further narrow the performance gaps, but it is not mandatory.
Caching
Query tokens do not deeply interact with each other. Thus, the query token vectors are cacheable. Given a certain vocabulary with size $V$, all token vectors $v_{t}^{den},t\in\{0,...,V-1\}$ are precomputed as in Equation 3 and stored in a lookup Embedding $E$.
Online serving
LLM-based encoders are no longer needed during online serving. Instead, only one Embedding layer, with a single weight matrix shaped $[V,H]$, is hosted in RAM. $H$ is the dimension of LLM’s hidden states, which is also the dimension of dense vectors. The dense query vector is obtained with a simple lookup-then-average operation,
$v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{den}=\frac{1}{n}\sum_{i=0}^{n-1}E[t_{i}]$ (6)
The online serving of our dense retriever is extremely lightweight, because such Embedding lookup is super efficient with or without GPU hardware.
2.3 Sparse retrieval of LightRetriever
The sparse retrieval of LightRetriever steps towards lightweight even further, by completely removing the need for the sparse query encoder.
Figure 4: Sparse retrieval of LightRetriever.
Training
The sparse query vector $v_{q}^{spr}$ directly maps a token and its corresponding quantity,
$v_{q}^{\text{spr}}[t]=\left\{\begin{array}[]{ll}0&\text{if }t\notin T_{q}\\ \text{Number of }t&\text{if }t\in T_{q}\end{array}\right.$ (7)
where $t\in\{0,...,V-1\}$ represents valid token ids within LLM’s vocabulary. We obtain the sparse document vector $v_{d}^{\text{spr}}$ by first feeding the document $d$ into the pretrained language model (LM), resulting in the final-layer hidden states $h_{\text{last}}=LM(d)$. These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$. Following SPLADE [20], to induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension,
$v_{d}^{\text{spr}}=\max(w)\text{, where }w=\text{logits}=\ln(\max(h_{\text{last}}\cdot P,0)+1)$ (8)
Here, $v_{d}^{\text{spr}}\in\mathbb{R}^{V}$ denotes the final sparse representation vector of the document. Given sparse vectors $v_{q}^{\text{spr}}$ and $v_{d}^{\text{spr}}$, the same contrastive loss in Equation 2 and auxiliary loss in Equation 5 are applied for representation learning.
$v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions. Thus, the sparsify regulator is needed during the training phase. Following SPLADE [20], given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator [56] is used for sparsification. This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
$\ell_{\text{FLOPS}}=\sum_{t=0}^{V-1}\left(\frac{1}{N}\sum_{i=0}^{N-1}w_{t}^{(d_{i})}\right)^{2}$ (9)
Online serving
The sparse query vector $v_{q}^{\text{spr}}$ during online serving follows the same rule in Equation 7. No more query encoder for sparse retrieval is needed.
3 Experiments
3.1 Settings
Datasets
A total of 23 training sets (including 20 English and 3 Chinese) are used in our experiments by following [47, 71]. Most of them are available at preprocessed data  [^4] https://huggingface.co/datasets/tdro-llm/finetune_data from tDRO [47] and Sentence Transformers Training Data  [^5] https://huggingface.co/datasets/sentence-transformers/embedding-training-data [59]. Detailed training sets and sampling ratios are listed in the Appendix A.3. BeIR [65] and CMTEB Retrieval [76] benchmarks are used in the evaluation. BeIR is a massive English retrieval benchmark collection with 15 evaluation sets. These heterogeneous datasets cover different domains, tasks, and retrieval granularities. CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains. nDCG@10 is reported as the main metric by following [65, 76], accessing both recall and ranking abilities. Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
Models & hyper-parameters
We modify the experiment code based on the implementation  [^6] https://github.com/ma787639046/tdro from tDRO [47]. All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps. Following SPLADE [20], the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to maximum at first 4k steps, which helps reduce the side effects of the regulator at initial steps [56]. LoRA [24] is used to save GPU memory usage during the LLM training, where LoRA r is 16, alpha is 32, and dropout is 0.1. Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Mistral-v0.3-7B [26], Llama-3.2-1B, 3B, Llama-3.1-8B [67], and Qwen-2.5-1.5B, 3B, 7B [2]. Direct linear interpolation [72] is used to combine dense and sparse similarity scores as the hybrid retrieval scores. All results are conducted on a DGX node with 8 NVIDIA H800 GPUs, 128 CPU cores of 2 Intel Xeon 8462Y+, and 2 TB of memory.
Baselines
Full-sized symmetric dense retrievers are our main baselines under the same training conditions, as shown in Figure 2. The query inference speeds and retrieval performance gaps are compared as the main results. Additionally, we also include some recent retrievers in the results, including Sentence Transformers static embedding [1], BM25 [60], BGE-m3 [11], E5-Mistral [71], and LLM2Vec [5]. However, they are not our main comparison baselines, because [1] and [60] are parameter-efficient but less performant. LLM-based retrievers [71, 5] are performant but inefficient.
Table 1: Comparison of query encoding time and speedup for full-sized baseline and LightRetriever. Tokenization time is also presented for reference. 65536 queries from MSMARCO [52] are processed. Evaluations are conducted on an NVIDIA DGX node using 1 H800 GPU with Mistral-v0.3-7B.
Tokenzation Baseline LightRetriever
GPU Encode No-GPU Encode GPU Encode
Batch Size Time (s) Time (s) Time (s) speedup Time (s) speedup
16 5.1168 69.9276 3.7065 19 0.1509 463
64 3.8995 42.2991 3.1719 13 0.0510 829
256 2.0713 45.8708 2.9990 15 0.0185 2480
1024 1.4226 57.4545 2.7969 21 0.0107 5370
4096 1.6798 74.7221 2.8218 26 0.0093 8035
8192 OOM 2.8529 0.0090
3.2 Main results
Encoding speeds
The encoding speed comparisons of full-sized baseline and LightRetriever are presented in Table 1. No-GPU and GPU settings are measured for LightRetriever. Because of the efficiency issue, the baseline is only measured on the GPU. As is shown in Table 1, it takes a full-sized Mistral-7B LLM up to 74s for encoding 65536 queries on an H800 GPU across different batch sizes. In comparison, our LightRetriever only consumes less than 0.1509s on the GPU, which achieves over 400-8000x speedup. Even when encoding without GPU acceleration, LightRetriever also achieves over 10-20x speedup compared to the GPU baseline.
Table 2: Retrieval performances of LightRetriever on BeIR and CMTEB Retrieval (CMTEB-R) benchmarks. The nDCG@10 gaps between LightRetriever hybrid and the corresponding baseline are presented. The best metrics of baselines and LightRetrievers are marked in bold.
Benchmark BEIR (15 datasets) CMTEB-R (8 datasets)
nDCG@10 R@20 R@50 R@100 nDCG@10 R@20 R@50 R@100
Static Embed 34.1 44.4 53.1 59.5 31.3 45.5 54.3 60.4
BM25 41.7 48.8 56.5 61.8 50.8 63.9 70.0 74.1
[…]
E5-Mistral 56.9 62.1 68.9 73.6 61.8 75.1 81.6 87.1
Llama3.2-1b
Baseline 53.1 60.2 66.9 72.0 63.8 78.1 84.1 87.9
Light Dense 48.7 55.8 63.0 68.3 58.7 73.1 79.7 83.8
Light Sparse 47.6 54.5 61.5 66.7 53.9 68.8 75.7 80.1
Light Hybrid 52.0-${}_{\text{-1.1}}$ 58.1 65.1 70.1 60.8-${}_{\text{-3.0}}$ 74.8 81.2 85.0
Llama3.2-3b
Baseline 55.6 62.9 69.7 74.7 66.1 81.0 86.6 90.0
Light Dense 51.0 58.6 65.5 70.5 60.4 75.1 81.5 85.3
Light Sparse 48.4 55.5 62.7 67.7 53.0 68.0 75.0 79.6
Light Hybrid 53.5-${}_{\text{-2.1}}$ 59.9 66.8 71.7 61.7-${}_{\text{-4.4}}$ 76.3 82.3 86.3
Llama3.1-8b
Baseline 56.8 64.3 71.1 75.7 67.6 82.0 87.4 90.9
Light Dense 52.0 59.5 66.5 71.4 61.4 75.6 82.3 86.4
Light Sparse 49.5 56.9 64.1 69.1 54.8 69.8 77.3 81.8
Light Hybrid 54.4-${}_{\text{-2.4}}$ 60.9 67.7 72.8 63.0-${}_{\text{-4.6}}$ 77.3 83.6 87.4
Mistral0.3-7b
Baseline 57.5 64.1 71.0 75.7 65.9 80.3 86.5 90.2
Light Dense 52.7 59.2 66.5 71.3 57.4 72.1 79.0 83.6
Light Sparse 48.0 56.1 63.3 68.3 50.7 65.5 72.7 77.8
Light Hybrid 54.5-${}_{\text{-3.0}}$ 60.1 67.4 72.1 59.4-${}_{\text{-6.5}}$ 73.9 80.6 84.8
Qwen2.5-1.5b
Baseline 54.0 60.5 67.5 72.3 65.9 81.0 86.7 90.4
Light Dense 48.9 56.0 63.1 68.1 61.9 77.0 83.2 87.4
Light Sparse 47.3 54.3 61.4 66.4 52.5 67.4 74.9 79.9
Light Hybrid 52.1-${}_{\text{-1.9}}$ 58.2 65.1 70.0 63.8-${}_{\text{-2.1}}$ 78.3 84.5 88.3
Qwen2.5-3b
Baseline 54.9 62.0 68.7 73.5 69.4 83.8 88.9 91.8
Light Dense 50.1 57.2 64.2 69.3 64.7 79.4 85.2 88.8
Light Sparse 47.5 54.9 62.0 67.2 54.1 69.2 76.5 81.1
Light Hybrid 52.8-${}_{\text{-2.1}}$ 59.1 66.0 70.9 65.7-${}_{\text{-3.7}}$ 80.2 86.1 89.3
Qwen2.5-7b
Baseline 56.6 63.6 70.5 75.2 70.1 84.4 89.4 92.4
Light Dense 50.6 58.6 65.6 70.5 65.0 79.8 85.8 89.3
Light Sparse 48.2 56.0 63.5 68.5 55.0 69.7 77.0 82.1
Light Hybrid 53.8-${}_{\text{-2.8}}$ 60.3 67.5 72.5 66.5-${}_{\text{-3.6}}$ 81.1 86.7 89.7
Retrieval performances
The retrieval performances are presented in Table 2. Experiments are conducted with 3 series of LLMs with different sizes. For each base model, we test the full-sized baseline and dense/sparse/hybrid (dense + sparse) retrieval performances of LightRetriever (Light). As is shown in the results, the performance gaps between LightRetriever (Hybrid) and its corresponding baseline are averaged around 5%, ranging from 2.1% (-1.1 point) to 9.9% (-6.5 point). Mistral [26] gets the highest nDCG@10 scores on the English BeIR benchmark, while Llama models [67] are better at recalls of BeIR. Qwen series [2] performs the best on the Chinese CMTEB Retrieval benchmark, attributed to better Chinese pre-training of Qwen.
Figure 5: Performances (Llama-3.1-8b) with different Top-k dimensions or sparsities.
3.3 Controls of vector dimension and sparsity
Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors. For example, indexing 8.8M MS-MARCO passages [52] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval. Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors. 1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) [35] with LightRetriever to support flexible controls of dense vector dimensions. Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors. MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k. 2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension. Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
We conduct experiments of controlling dense vector dimension and sparse vector sparsity with Top-k of $\{128,256,512,1024,2048,4096\}$ on Llama-3.1-8b. As presented in Figure 5, the performances of LightRetriever decrease marginally when shrinking Top-k to 4x ($4090\rightarrow 1024$). The lowest Top-k setting ($128$) has 7.9% and 8.8% performance gaps on BeIR and CMTEB-R benchmarks compared to the highest Top-k ($4096$), but its disk usages with MS-MARCO are only 4.22 GB for dense trieval and 3.03 GB for sparse retrieval. This shows that our LightRetriever is capable of flexible index controls via simple vector shrinking techniques, with limited effect on retrieval performances.
Table 3: Ablations (nDCG@10, w/Llama-3.1-8b) on BeIR and CMTEB-R about auxiliary KL loss.
Benchmark w/ KL wo/ KL
BeIR CMTEB-R BeIR CMTEB-R
Baseline 56.8 67.6 57.0+0.2 67.9+0.3
Light Dense 52.0 61.4 50.8-1.2 60.1-1.3
Light Sparse 49.5 54.8 48.7-0.8 54.4-0.4
Light Hybrid 54.4 63.0 53.9-0.5 62.1-0.9
3.4 Effect of auxiliary KL loss
Previous works for BERT-based asymmetric retrievers [73, 32] emphasize the importance of alignment between asymmetric students and full symmetric teachers. Our work also uses the KL loss as an auxiliary. Results in Table 3 show that removing such loss incurs limited drops of -0.5 and -0.9 on BeIR and CMTEB-R. Thus, such auxiliary KL loss is not mandatory in our experiment settings.
Table 4: Ablations of dense and sparse retrieval for LightRetriever (Llama-3.2-1b).
Dense Ablation BeIR CMTEB-R
LightRetriever 48.7 58.7
Symmetry
w/ Qry Full - Doc Light 8.7-40.0 9.4-49.3
w/ Qry Light - Doc Light 34.9-13.8 40.1-18.6
Other Ablations
w/ Direct Emb Layer 37.5-11.2 41.3-17.4
wo/ Instruction 47.1-1.6 56.1-2.6
Sparse Ablation BeIR CMTEB-R
LightRetriever 47.6 53.9
Symmetry
w/ Qry Full - Doc Light 16.2-31.4 10.8-43.1
Other Ablations
w/ Value fixed to 1 47.6+0.0 54.7+0.8
4 Ablations
4.1 Symmetry
Our work proposes an asymmetric architecture with an extremely lightweight query (Qry) encoder and a full-sized document (Doc) encoder. Besides, we also perform ablations on other symmetry settings, including lightweight on both query and document, and a full-sized query encoder with a lightweight passage encoder. Results in Table 4 show that a full-sized query encoder with a lightweight passage encoder leads to drastic performance degenerations, maybe caused by embedding collapses. Lightweight on both sides for dense retrieval works, but it does not perform excellently in retrieval tasks, just like Static Embedding [1]. Lightweight on both sides is not suitable for our sparse retrieval, because no learnable parameters are available.
4.2 Other ablations
Do we need a full query model during the dense training?
Dense retrieval of LightRetriever trains token vectors with full Transformers. Fast retrieval is achieved by caching learned dense token vectors and lookups during serving. If we directly use the Embedding Layer within the Transformers, the performances drop 11.2 and 17.4 points on BeIR and CMTEB-R. This verifies that during the training phase, a full-sized model is still needed for learning query token vectors.
Task instructions
Task instructions contribute +1.6 and +2.6 points on BeIR and CMTEB-R of dense retrieval. Similar effects of instructions are also observed in previous LLM retrievers [71, 5].
Value of sparse query vector
The sparse query vector is a key-value type of mapping. In Equation 7, we model the value of sparse query vectors by counting the corresponding token ids that appear in the query. For ablations, we choose a simpler way of modeling by stopping the counting and fixing the value to 1. In Table 4, BeIR results do not change. However, CMTEB-R results improve by +0.8 points. This indicates that counting term frequency in our sparse query vector seems unimportant to some extent, under large amounts of training on LightRetriever. The co-occurrence is more important.
5 Conclusion
Existing LLM-based retrievers use symmetric dual-encoders to model both queries and documents. While documents could be pre-encoded and indexed offline, queries arrive in real-time and need online encoding. The deployment of large LLMs results in inefficient online services. In this paper, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture capable of extremely lightweight query encoding. Our approach achieves over a 1000x speedup for query inference on an H800 GPU, even a 20x speedup without GPU acceleration. Experiments show great robustness and generalization abilities across different foundation LLMs and retrieval tasks.
Appendix A Technical Appendices and Supplementary Material
A.1 Limitations
Our work focuses on online efficiency optimizations of recent LLM-based architecture, whose giant parameters pose a greater need for such optimizations. Other readers may be interested in the application of our methods on non-LLM architectures, e.g., BERT. While such a combination is possible, due to limited time constraints and lesser needs for non-LLM, we leave this to future work.
A.2 Related works
A.2.1 Text retrieval
Text retrieval builds on top of the vector space model [62], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
Dense retrieval
Dense retrieval [29] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) [51] algorithm. In recent years, dense retrievers, such as BGE [11] and E5 [70], have gained popularity for strong retrieval abilities across different tasks [50, 65], languages [83], and context granulaties [74, 52], owing to diverse training data [59], retrieval-customized pre-training [21, 75, 46], improved negative-mining [78], and so on. Recent SOTA retrievers [71, 5, 38] start to utilize LLMs [26, 67] as backbone encoders. With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities. However, LLMs also incur heavy online workloads when serving as query encoders. Improving the serving efficiency of LLM-based retrievers is still left for exploration.
Sparse retrieval
Traditional statistical sparse retrieval algorithms (e.g., TF-IDF [63], and BM25 [60]) do not involve learnable language models. Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space. These lexicals are sparsely distributed and require building inverted indexes [45] for feasible similarity metric calculation, such as Apache Lucene [6] and Tantivy [58]. In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs). BGE [11], SPLADE [20], SparTerm [3], and SPARTA [84] use customized XLM-RoBERTa [14] or BERT [18] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors. Because of sparsity, these works also search with inverted indexes. Recent works [85] also try to incorporate LLMs with sparse retrieval. However, the online efficiency issue still exists.
Hybrid retreval
Hybrid retrieval is a technique that combines retrieval scores [72] or ranks [15] from multiple retrieval systems, such as dense [29] / sparse [60, 20] retrieval, and single [29] / multiple [31, 11] vector retrieval. It enables better retrieval abilities by interpolating results from multiple sources. Linear interpolation [72] of dense and sparse scores is used in our work.
A.2.2 Improving inference efficiency
Numerous efforts endeavor to improve the inference efficiency of text retrieval.
Distilling smaller encoders
Previous works have tried to distill small query encoders from large encoders for better query inference efficiency. KALE [9] prunes the BERT-based query encoder and distills it during post-training. However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ [36] decreases by 25-30%. [73] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models. Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization. USTAD [32] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students. Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
However, these studies have primarily focused on small language models (SLMs), such as BERT [18], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical. Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains. In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO [52], thus unable to compete with LLM-based retrievers.
In contrast, with the rapid advancement of LLM-based text retrievers, driven by larger and more diverse foundation models and datasets, retrieval performance has significantly improved. To address the growing need for efficient online inference in this setting, we propose a novel method that, to our knowledge, is the first to target the online efficiency challenge in LLM-based retrievers. Our approach eliminates the need for Transformers layers for query encoders and does not rely on mandatory distillation. Yet our work only incurs an average performance drop of about 5%. Moreover, we evaluate our method across multiple LLMs of varying sizes and diverse retrieval tasks, demonstrating strong generalization capabilities.
Shrinking index sizes
Previous works have also been dedicated to reducing index sizes for smaller disk space usage. Matryoshka representation learning (MRL) [35] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors. It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning. For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation. For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals [20]. Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals. To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Experiments Section 3.3, which also shows strong generalization capacities.
A.3 Training datasets
Dataset infos
Following the settings of previous LLM-based retrievers [71, 5, 47], our work uses 23 training sets pre-built by tDRO [47] or Sentence Transformers Training Data [59], including 20 English and 3 Chinese datasets. Detailed dataset information is listed in Table 5 as follows.
Table 5: Training Dataset informations.
Dataset Language Category Deduped Size Epoch Ratio
Amazon Review (2018) [53] English Amazon 999999 0.01 0.58%
cMedQA2 [81] Chinese Chinese Medical 99936 1 5.84%
DuReader [57] Chinese Chinese Web Collections 86366 1 5.05%
Eli5 [19] English Reddit 325390 0.05 0.95%
Fever [66] English Wikipedia QA 109808 0.8 5.13%
FiQA [48] English Financial 5498 1 0.32%
GooQA Pairs [30] English Web Collections 3012347 0.025 4.40%
HotpotQA [80] English Wikipedia QA 85000 0.5 2.48%
MSMARCO [52] English Web Collections 502854 1 29.39%
NFCorpus [7] English Medical 2585 1 0.15%
NPR [44] English News 594376 0.05 1.74%
NQ [36] English Wikipedia QA 58800 1 3.44%
PQA Pairs [40] English Wikipedia QA 99999 0.5 2.92%
Quora Duplicates Triples [25] English Forum Duplicates 97011 0.5 2.83%
S2ORC Title-Abstract [42] English Semantic Scholar 99998 0.5 2.92%
SciFact [69] English S2ORC 806 1 0.05%
SPECTER [13] English Semantic Scholar 136642 0.25 2.00%
StackExchange Duplicates (Title-Body) [16] English Forum Duplicates 250516 0.25 3.66%
E5 Synthetic [4] English GPT Synthetic Data 224791 1 13.14%
T2Ranking [77] Chinese Chinese Web Collections 200362 0.5 5.86%
Trivia [28] English Wikipedia QA 60370 0.5 1.76%
Wikihow [33] English WikiHow 128543 0.25 1.88%
Yahoo Answers (Title-Answer) [82] English Yahoo 1198018 0.05 3.50%
Many downstream retrieval benchmarks [65, 76] are evaluated out-of-domain, where no training samples from the downstream sources are available. To avoid the overfitting of training sets, as is shown in the above Table 5, we slightly tune the epoch used in the dataset sampling, by following previous works in [54, 71, 47].
Training set decontamination
Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems. Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets. Following tDRO [47], we perform strict training set decontamination with SimHash [49] to ensure no training queries appear in the BeIR [65] and CMTEB Retrieval [76]. The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 6.
Table 6: Training set deduplications. We report each dataset with the original size, deduplicated size, and number of duplicates detected in both the train and test sets.
Dataset Original Size Deduped Size Duplicates
Amazon Review (2018) 1000000 999999 1
cMedQA2 100000 99936 64
DuReader 86395 86366 29
Eli5 325475 325390 85
Fever 109810 109808 2
FiQA 5498 5498 0
GooQA Pairs 3012496 3012347 149
HotpotQA 85000 85000 0
MSMARCO 502939 502854 85
NFCorpus 2590 2585 5
NPR 594384 594376 8
NQ 58812 58800 12
PQA Pairs 100000 99999 1
Quora Duplicates Triples 101762 97011 4751
S2ORC Title-Abstract 100000 99998 2
SciFact 809 806 3
SPECTER 136645 136642 3
StackExchange Duplicates (Title-Body) 250519 250516 3
E5 Synthetic 224791 224791 0
T2Ranking 200376 200362 14
Trivia 60380 60370 10
Wikihow 128543 128543 0
Yahoo Answers (Title-Answer) 1198260 1198018 242
Additionally, we also listed some samples of the duplicated training queries, as well as the corresponding test sets in Table 7.
Table 7: Samples of duplicate queries found in both the train and test sets.
Train Set Train qid Train Query Test Set Test qid Test Query
Quora train-164 Is there a framework for auditing social media? Quora 201573 Is there a framework for auditing social media?
Yahoo-Answer train-10046 who is sachin tendulkar? Quora 47702 Who is Sachin Tendulkar?
MSMARCO 968274 where did abraham lincoln died DBPedia QALD2_tr-6 Where did Abraham Lincoln die?
Eli5 train-3088 How does unemployment insurance work? FiQA 2648 How does unemployment insurance work?
A.4 Task instructions
As previously mentioned in Section 2.1, modern LLM-based retrievers [71, 5] introduce task-specific instructions to format input queries, which enables higher retrieval capacities. These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation [55]. Following Mistral-E5 [71], we reuse most of its task instructions, and format the instruction and query in the following format:
[…]
The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral [71] and tDRO [47] for reproducibility. Note that one training set may correspond to multiple instructions to avoid overfitting.
Instructions for training sets
  1. 1.
    Amazon Review (2018)(1): Given a title, retrieve the corresponding reviews from Amazon
  2. 2.
    Amazon Review (2018)(2): Given a title, retrieve a Amazon review
  3. 3.
    cMedQA2: Given a Chinese community medical question, retrieve replies that best answer the question
  4. 4.
    DuReader: Given a Chinese search query, retrieve web passages that answer the question
  5. 5.
    Eli5: Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
  6. 6.
    GooQA Pairs: Given a web search query, retrieve the corresponding answers from Google
  7. 7.
    HotpotQA: Given a multi-hop question, retrieve documents that can help answer the question
  8. 8.
    MSMARCO: Given a web search query, retrieve relevant passages that answer the query
  9. 9.
    NQ(1): Given a question, retrieve Wikipedia passages that answer the question
  10. 10.
    NQ(2): Retrieve Wikipedia passages that answer the question
  11. 11.
    Quora Duplicates Triples(1): Given a question, retrieve questions that are semantically equivalent to the given question
  12. 12.
    Quora Duplicates Triples(2): Find questions that have the same meaning as the input question
  13. 13.
    S2ORC Title-Abstract(1): Given a title, retrieve the abstract from scientific papers
  14. 14.
    S2ORC Title-Abstract(2): Given a title, retrieve abstracts from scientific papers that match the title
  15. 15.
    SciFact: Given a scientific claim, judge whether the document supports or refutes the claim
  16. 16.
    SPECTER(1): Given a title, retrieve semantic related titles
  17. 17.
    SPECTER(2): Retrieve semantic related titles from scientific publications
  18. 18.
    StackExchange Duplicates (Title-Body): Retrieve duplicate questions and passages from StackOverflow forum
  19. 19.
    T2Ranking: Given a Chinese search query, retrieve web passages that answer the question
  20. 20.
    Trivia(1): Given a question, retrieve Wikipedia passages that answer the question
  21. 21.
    Trivia(2): Retrieve Wikipedia passages that answer the question
  22. 22.
    Wikihow: Given a summary, retrieve Wikipedia passages that match the summary
  23. 23.
    Yahoo Answers (Title-Answer): Given a title, retrieve Yahoo answers that match the title
Instruction for English BeIR benchmarks
  1. 1.
    ArguAna: Given a claim, find documents that refute the claim
  2. 2.
    ClimateFEVER: Given a claim about climate change, retrieve documents that support or refute the claim
  3. 3.
    CQADupStack: Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
  4. 4.
    DBPedia: Given a query, retrieve relevant entity descriptions from DBPedia
  5. 5.
    FEVER: Given a claim, retrieve documents that support or refute the claim
  6. 6.
    FiQA2018: Given a financial question, retrieve user replies that best answer the question
  7. 7.
    HotpotQA: Given a multi-hop question, retrieve documents that can help answer the question
  8. 8.
    MSMARCO: Given a web search query, retrieve relevant passages that answer the query
  9. 9.
    NFCorpus: Given a question, retrieve relevant documents that best answer the question
  10. 10.
    NQ: Given a question, retrieve Wikipedia passages that answer the question
  11. 11.
    Quora: Given a question, retrieve questions that are semantically equivalent to the given question
  12. 12.
    SCIDOCS: Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
  13. 13.
    SciFact: Given a scientific claim, retrieve documents that support or refute the claim
  14. 14.
    Touche2020: Given a question, retrieve detailed and persuasive arguments that answer the question
  15. 15.
    TRECCOVID: Given a query on COVID-19, retrieve documents that answer the query
Instruction for Chinese CMTEB Retrieval benchmarks
  1. 1.
    CmedqaRetrieval: Given a Chinese community medical question, retrieve replies that best answer the question
  2. 2.
    CovidRetrieval: Given a question on COVID-19, retrieve news articles that answer the question
  3. 3.
    DuRetrieval: Given a Chinese search query, retrieve web passages that answer the question
  4. 4.
    EcomRetrieval: Given a user query from an e-commerce website, retrieve description sentences of relevant products
  5. 5.
    MMarcoRetrieval: Given a web search query, retrieve relevant passages that answer the query
  6. 6.
    MedicalRetrieval: Given a medical question, retrieve user replies that best answer the question
  7. 7.
    T2Retrieval: Given a Chinese search query, retrieve web passages that answer the question
  8. 8.
    VideoRetrieval: Given a video search query, retrieve the titles of relevant videos
A.5 Efficiency optimizations techniques
Customized causal mask
As defined in the Equation 3, our work computes dense token vectors for each input token with a common prompt. This causes the repeated computation of the common prompt area. We develop a customized causal mask for LLMs to avoid such computation waste. Specifically, given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, we format the instruction and query tokens as follows,
[…]
Figure 6: An example of the customized attention mask.
The customized causal mask ensures that the token blocks can attend to the common prompt area, but not attend to each other. Let the sequence length be $L$, prompt length be $P$, micro block (each token + <eos>) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index. The attention mask (0 means attend, $\infty$- ∞ means masked) is defined as,
$\text{Mask}[q,k]=\begin{cases}0,&\text{if }q<P\text{ and }k\leq q\quad\text{(attn within prompt)}\\ 0,&\text{if }q\geq P\text{ and }k<P\quad\text{(blocks can attend to prompt)}\\ 0,&\text{if }q\geq P\text{ and }\left\lfloor\frac{q-P}{w}\right\rfloor=\left\lfloor\frac{k-P}{w}\right\rfloor\text{ and }k\leq q\quad\text{(attn within current block)}\\ -\infty,&\text{otherwise}\end{cases}$
An example of the customized causal mask is shown in Figure 6.
Optimizations of sparse max aggregation
As stated in Equation 8, the sparse vector is computed via
[…]
, where the vocabulary dimension $V$ is much larger than the hidden dimension $H$. For example, for Llama3.1-8b [67] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$. This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference. However, only a tensor with shape $[B,V]$ is needed for the final aggregated representations. Thus, an optimization for computing the sparse vector is needed urgently.
We design an efficient way of computing the sparse vector. Firstly, because ReLU and Log are monotonically increasing functions, we first move the last max aggregation operations to step 2, right after the projection.
[…]
Then we design a customized PyTorch Autograd function to fuse steps 1&2. It computes the projection and maximum operations together along the sequence length dimension $T$. Such sliced computation avoids the creation of a large tensor with a sequence length dimension. Our preliminary observation shows that this helps save the GPU memory usage of the sparse max aggregation up to 100-200x. The forward and backward are presented in the Algorithm 1.
Algorithm 1 Forward and Backward Pass of Efficient Projection then Max aggregation
1:Input tensor $X\in\mathbb{R}^{B\times T\times H}$, weight $W\in\mathbb{R}^{H\times V}$, optional bias $b\in\mathbb{R}^{V}$, optional input attention mask $M\in\{0,1\}^{B\times T}$, where B is batch size, T is sequence length, H is hidden size, V is vocabulary size.
2:Output tensor $Y\in\mathbb{R}^{B\times V}$
3:function Forward($X,W,b,M$)
4:     Initialize $Y\leftarrow-\infty$
5:     for $t=0$ to $T-1$ do
6:         $L_{t}\leftarrow X[:,t,:]\cdot W$
7:         if $b$ is not None then
8:              $L_{t}\leftarrow L_{t}+b$
9:         end if
10:         if $M$ is not None then
11:              for $i=0$ to $B-1$ do
12:                  if $M[i,t]=0$ then
13:                       $L_{t}[i,:]\leftarrow-\infty$
14:                  end if
15:              end for
16:         end if
17:         $Y\leftarrow\max(Y,L_{t})$
18:     end for
19:     Save $\text{argmax}_{t}$ indices as max_indices
20:     return $Y$
21:end function
22:function Backward($\nabla Y,\texttt{max\_indices},X,W$)
23:     Initialize $\nabla X\leftarrow 0$, $\nabla W\leftarrow 0$, $\nabla b\leftarrow 0$
24:     for $t=0$ to $T{-}1$ do
25:         $X_{t}\leftarrow X[:,t,:]\in\mathbb{R}^{B\times H}$
26:         $\nabla X_{t}\leftarrow\nabla X[:,t,:]\in\mathbb{R}^{B\times H}$
27:         $M_{t}\leftarrow\texttt{max\_indices}_{I==t}\in\{0,1\}^{B\times V}$
28:         $\nabla L_{t}\leftarrow\nabla Y\odot M_{t}\in\mathbb{R}^{B\times V}$triangleright Only keep grads where $t$ was max
29:         $\nabla X_{t}\mathrel{+}=\nabla L_{t}\cdot W^{\top}$
30:         $\nabla W\mathrel{+}=X_{t}^{\top}\cdot\nabla L_{t}$
31:         $\nabla b\mathrel{+}=\sum_{i=1}^{B}\nabla L_{t}[i,:]$
32:     end for
33:     return $\nabla X,\nabla W,\nabla b$
34:end function
Other important optimizations
Proper optimizations are important. For example, we found it hard to reproduce quickly with some of the open-sourced baselines [5], which is roughly 6x slower than our implementation when testing with the same size of LLM. Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently. We import other optimizations based on the implementations  [^7] Please refer to [47]’s implementation on https://github.com/ma787639046/tdro. released by tDRO [47]. These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC [41]; 2) Sequence packing [34] with Flash Attention 2 [17], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel [23], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache [22] for optionally enlarging batch sizes if necassary.
A.6 Detailed main results
The detailed main results (nDCG@10) on each dataset are shown in the tables below, which are full versions of Table 2.
Table 8: (Part1) Detailed main results (nDCG@10) on BeIR test sets (except MSMARCO, it uses the dev set).
Task / Test, nDCG@10 ArguAna CQADup CFEVER DBPedia FEVER FiQA HotpotQA MSMARCO
Static Embedding 44.4 22.0 20.4 27.4 43.0 20.0 46.8 17.9
BM25 31.5 29.9 21.3 31.3 75.3 23.6 60.3 22.8
BGE-m3-dense 53.9 38.4 26.5 38.5 77.2 40.8 68.6 37.6
BGE-m3-sparse 36.2 30.5 26.7 24.4 85.9 26.9 70.3 17.1
LLM2Vec-llama3-8b 62.8 48.3 34.3 48.3 90.2 55.3 71.8 43.2
E5-Mistral 61.9 42.9 38.3 48.9 87.8 56.6 75.7 43.1
Llama3.2-1b
Baseline 58.0 44.0 41.2 43.2 89.2 45.5 69.3 42.7
Light Dense 49.2 38.1 37.9 34.2 83.6 40.4 60.8 39.2
Light Sparse 49.5 34.5 27.5 35.9 84.7 35.7 65.5 37.2
Light Hybrid 54.1 39.7 39.3 39.5 87.3 42.1 68.2 41.1
Llama3.2-3b
Baseline 58.5 46.8 43.7 46.4 89.9 53.4 73.7 44.3
Light Dense 55.3 40.3 41.7 36.5 84.0 46.2 64.3 41.6
Light Sparse 46.5 34.1 26.9 38.5 86.2 38.9 66.3 39.2
Light Hybrid 56.3 41.4 42.4 41.6 87.6 46.9 69.7 42.9
Llama3.1-8b
Baseline 57.1 49.2 44.1 48.2 90.9 59.0 77.8 46.2
Light Dense 55.8 40.9 41.5 38.7 85.7 49.8 67.6 43.4
Light Sparse 46.4 33.6 32.2 38.8 86.1 41.9 67.9 41.0
Light Hybrid 57.2 41.9 42.4 42.0 87.8 50.7 71.5 44.5
Mistral0.3-7b
Baseline 56.2 49.7 43.6 48.4 91.3 59.6 77.7 46.6
Light Dense 53.5 41.6 40.9 38.9 85.6 50.9 66.9 44.2
Light Sparse 44.0 33.6 21.9 37.0 81.0 43.6 65.4 40.9
Light Hybrid 54.3 42.2 41.4 40.5 86.9 51.7 69.9 44.8
Qwen2.5-1.5b
Baseline 56.6 44.1 40.4 43.9 88.9 47.9 68.1 42.6
Light Dense 51.2 38.3 36.1 33.0 82.1 41.2 59.0 39.1
Light Sparse 46.5 35.0 25.7 36.3 85.1 35.3 63.3 37.3
Light Hybrid 54.1 40.1 37.7 39.5 87.0 42.1 66.5 40.8
Qwen2.5-3b
Baseline 55.5 47.0 44.1 45.8 89.6 52.9 71.6 43.7
Light Dense 49.8 39.6 41.1 35.0 83.9 45.4 62.8 41.0
Light Sparse 45.5 36.2 28.7 36.9 84.9 38.6 64.2 38.7
Light Hybrid 52.8 41.4 42.1 39.8 86.9 46.1 68.2 42.3
Qwen2.5-7b
Baseline 57.1 48.2 43.5 47.9 90.7 56.5 74.7 44.8
Light Dense 53.3 40.0 39.1 36.8 85.1 48.8 64.4 41.8
Light Sparse 47.5 35.6 28.2 36.7 85.4 40.2 65.2 40.1
Light Hybrid 54.4 41.6 42.2 41.0 87.9 49.7 69.4 43.3
Table 9: (Part2) Detailed main results (nDCG@10) on BeIR test sets.
Task / Test, nDCG@10 NFCorpus NQ Quora SCIDOCS SciFact TRECCOVID Touche2020 Avg (# 15)
Static Embedding 30.0 23.1 77.5 13.2 59.3 44.6 22.4 34.1
BM25 32.5 32.9 78.9 15.8 66.5 65.6 36.7 41.7
BGE-m3-dense 31.5 60.1 88.4 15.2 63.6 47.0 20.6 47.2
BGE-m3-sparse 28.3 20.3 73.4 12.2 63.3 52.7 27.8 39.7
LLM2Vec-llama3-8b 41.8 64.2 87.2 23.0 78.2 80.3 20.5 56.6
E5-Mistral 38.6 63.5 89.6 16.3 76.4 87.2 26.4 56.9
Llama3.2-1b
Baseline 35.5 58.6 89.4 19.8 72.3 66.7 20.9 53.1
Light Dense 30.1 52.2 87.0 17.2 66.5 68.8 24.7 48.7
Light Sparse 34.0 52.0 82.2 17.2 69.1 60.6 27.8 47.6
Light Hybrid 34.2 56.6 87.4 18.6 71.5 70.5 30.2 52.0
Llama3.2-3b
Baseline 37.6 62.8 89.5 22.5 75.3 68.7 21.2 55.6
Light Dense 31.7 56.4 87.8 19.0 70.4 68.6 21.6 51.0
Light Sparse 33.3 55.3 83.2 17.7 67.3 64.1 28.5 48.4
Light Hybrid 34.2 59.8 88.1 20.0 71.4 70.9 29.5 53.5
Llama3.1-8b
Baseline 38.6 65.5 89.9 23.8 77.3 62.0 22.2 56.8
Light Dense 31.8 59.0 88.5 19.3 70.6 65.7 21.1 52.0
Light Sparse 33.6 58.4 84.4 18.9 69.7 63.0 27.1 49.5
Light Hybrid 34.6 61.9 88.6 20.7 72.6 71.5 27.8 54.4
Mistral0.3-7b
Baseline 36.8 66.9 90.0 22.7 77.7 71.3 23.6 57.5
Light Dense 31.9 59.8 88.7 18.5 70.3 71.5 26.7 52.7
Light Sparse 33.1 57.9 84.6 18.1 66.7 61.1 31.7 48.0
Light Hybrid 33.8 62.5 88.8 19.8 71.6 76.7 32.4 54.5
Qwen2.5-1.5b
Baseline 36.4 57.1 89.0 20.4 72.7 73.0 28.4 54.0
Light Dense 29.8 50.9 86.7 18.1 66.5 72.6 29.1 48.9
Light Sparse 32.9 51.6 81.3 17.4 66.4 64.8 31.1 47.3
Light Hybrid 34.0 55.8 87.0 19.3 69.2 76.8 32.3 52.1
Qwen2.5-3b
Baseline 35.9 61.4 89.5 22.3 72.9 65.7 26.0 54.9
Light Dense 30.6 54.5 87.5 18.5 64.8 71.2 26.1 50.1
Light Sparse 33.2 54.2 82.2 17.6 67.6 57.0 26.4 47.5
Light Hybrid 34.2 58.5 87.8 19.5 69.1 74.2 29.8 52.8
Qwen2.5-7b
Baseline 37.2 63.7 89.7 23.7 77.2 69.1 25.3 56.6
Light Dense 30.7 55.9 87.6 19.5 67.9 64.1 24.1 50.6
Light Sparse 33.9 55.5 78.7 18.3 69.0 60.5 28.5 48.2
Light Hybrid 34.4 60.1 87.9 21.1 71.1 73.1 29.4 53.8
Table 10: Detailed main results (nDCG@10) on CMTEB Retrieval dev sets.
Task / Dev, nDCG@10 Cmedqa Covid DuReader Ecom MMarco Medical T2 Video Avg (# 8)
Static Embedding 7.5 45.7 37.9 33.9 34.9 15.7 35.8 38.9 31.3
BM25 13.7 86.6 57.1 45.1 48.3 32.1 60.5 62.7 50.8
BGE-m3-dense 31.0 77.2 82.9 57.5 76.9 51.5 80.8 54.4 64.0
BGE-m3-sparse 24.5 76.0 71.4 50.3 59.2 44.0 71.7 58.5 57.0
LLM2Vec-llama3-8b 35.2 16.5 80.7 54.4 76.5 54.6 65.2 52.4 54.4
E5-Mistral 34.2 73.1 87.0 46.0 74.8 52.8 80.7 45.4 61.8
Llama3.2-1b
Baseline 32.3 74.2 84.6 53.8 73.5 50.5 78.0 63.1 63.8
Light Dense 27.2 67.6 80.8 51.0 67.9 43.0 73.4 58.5 58.7
Light Sparse 17.3 64.3 74.4 51.0 63.3 39.4 69.3 52.2 53.9
Light Hybrid 27.2 69.7 81.8 55.5 69.8 44.1 75.9 62.3 60.8
Llama3.2-3b
Baseline 36.8 75.6 87.9 57.7 75.7 53.8 81.1 60.5 66.1
Light Dense 29.8 70.2 84.0 51.7 69.2 46.5 75.7 56.4 60.4
Light Sparse 17.2 65.4 76.1 47.4 65.0 42.5 68.5 42.0 53.0
Light Hybrid 29.9 71.5 84.8 54.5 71.2 47.2 77.1 57.7 61.7
Llama3.1-8b
Baseline 39.5 77.1 89.5 58.9 78.1 57.6 84.1 56.2 67.6
Light Dense 32.1 71.0 85.7 53.0 70.2 49.0 78.5 52.0 61.4
Light Sparse 15.6 70.4 82.9 47.3 63.9 42.2 75.1 41.3 54.8
Light Hybrid 32.1 72.8 86.6 55.3 71.6 49.7 80.1 55.8 63.0
Mistral0.3-7b
Baseline 38.5 74.1 88.6 54.7 75.6 56.6 82.3 56.5 65.9
Light Dense 30.4 64.4 83.6 47.0 64.3 46.3 74.7 48.5 57.4
Light Sparse 11.5 63.2 80.5 43.4 61.7 36.7 71.3 37.4 50.7
Light Hybrid 30.3 67.2 84.5 50.7 67.6 46.9 76.4 51.4 59.4
Qwen2.5-1.5b
Baseline 37.9 78.2 87.6 56.9 75.9 54.7 78.8 57.0 65.9
Light Dense 31.3 73.8 85.3 52.8 71.8 49.5 75.2 55.7 61.9
Light Sparse 14.5 66.7 78.1 49.3 61.6 39.1 63.5 47.0 52.5
Light Hybrid 31.2 74.4 86.0 56.5 73.4 50.5 76.4 61.6 63.8
Qwen2.5-3b
Baseline 39.4 80.1 89.5 61.9 77.8 58.7 84.4 63.6 69.4
Light Dense 32.7 76.0 86.8 56.7 73.2 51.6 80.9 59.5 64.7
Light Sparse 12.8 70.4 80.2 48.8 61.1 39.4 72.4 47.9 54.1
Light Hybrid 32.6 77.0 87.6 58.3 74.0 51.6 82.0 62.8 65.7
Qwen2.5-7b
Baseline 41.3 81.4 91.0 61.1 80.0 59.8 85.5 60.3 70.1
Light Dense 34.3 74.6 88.0 55.7 75.4 53.2 81.6 57.2 65.0
Light Sparse 12.5 72.5 82.4 48.3 65.2 40.4 73.3 45.2 55.0
Light Hybrid 34.0 76.7 88.6 58.0 76.2 53.9 83.0 61.4 66.5